OpenStudy (anonymous):

What are the x– and y–intercepts of y = 2x^2 – 10x + 8?

OpenStudy (anonymous):

x intercepts take y=0

OpenStudy (anonymous):

0=2x^2 – 10x + 8

OpenStudy (anonymous):

solve for x

OpenStudy (anonymous):

(factor)

OpenStudy (anonymous):

y intercepts make x=0 so y=8

OpenStudy (merchandize):

To find your y intercept, plug 0 in for x. 2(0)^2 - 10*0 +8 y=8 To find your x intercepts, factor. You can pull a 2 out of all of the numbers. 2(x^2 - 5x +4) This makes factoring easier. 2(x-4)(x-1) Now, put each part with an x equal to zero. x-4=0 x=4 x-1=0 x=1 Therefore, your x intercepts are 4 and 1 and your y intercept is 8.

OpenStudy (anonymous):

thanks i get it now